# Dynamics of feedback Ising model

### v1

@@ III.1 Zero external magnetic field
When temperature approaches to zero, $\beta\to\infty$, the $M_{+}$ phase approaches to the plus ground state. A more precise asymptotic
@@ III.2 Nonzero external magnetic field
when $\gamma\to 0$. In the complimentary $\gamma\to+\infty$ limit
@@ III.3 Maxwell curve
for the FIM, while the ‘potential energy’ is
$\begin{split}V(m)&=\int_{0}^{m}du\,\text{arctanh}(u)\\ &=m\cdot\text{arctanh}(m)+\frac{1}{2}\log(1-m^{2})\\ &=\frac{1}{2}\Big[(1+m)\log(1+m)+(1-m)\log(1-m)\Big].\end{split}$ (61)
@@
Although a unique Maxwell magnetic field exists at a fixed temperature, a unique Maxwell temperature may not exist at a fixed magnetic field. In the classical IM with $\gamma=0$, we have $h_{*}(T)=0$ as shown in Fig. 3, so there are infinitely many Maxwell temperatures at $h=0$ and none otherwise. Thus, increasing the temperature favors neither the upper nor the lower equilibrium. However, in the FIM with $\gamma>0$, $h_{*}(T)$ increases to the cusp $(T_{c},h_{c})$ as shown in Figs. 33, so there is a unique Maxwell temperature for any fixed $h\in(h_{*}(0),h_{c})$ and none otherwise. For any fixed $h\in(0,h_{c})$, the Maxwell temperature exists in the temperature island disconnected from zero temperature. The unique feature of the FIM is that the Maxwell temperature is well defined when $h$ is small and either positive or negative, and increasing the temperature favors the lower equilibrium.
@@ III.5 Convergence to stable equilibrium
Figure 10: The evolution of the probability $P_{M_{1}}(t)$ of being in the adsorbing state. We chose $M_{0}=2M_{1}+1$ as the initial magnetization. In this case, the adsorption probability is (67). For the three shown curves, $M_{1}+1=30,100,300$. All curves intersects at $t=\log 2$ where $P_{M_{1}}=\frac{1}{2}$. The front around $t=\log 2$ becomes sharper as $M_{1}$ increases.
@@ IV.1 Fokker-Planck equation and approximation error
Hereinafter we use the variable $m\equiv 2x-1$, and we shortly write $\rho$ instead of $\rho(m,t)$. The FP equation in this case reads
@@
$\frac{\log(r)}{2}-\frac{r-1}{r+1}=\frac{(r-1)^{3}}{24}-\frac{(r-1)^{4}}{12}+\ldots$ (78)
@@ IV.2 Equilibrium distributions near critical points
$\alpha=\frac{\partial a}{\partial x},\quad\beta_{1}=\frac{1}{2}\frac{\partial^{2}a}{\partial y^{2}},\quad\beta_{2}=\frac{\partial^{2}a}{\partial m\partial y},\quad\sigma=\frac{1}{2}\frac{\partial^{2}a}{\partial m^{2}},$ (84)
@@
$\begin{split}&-\epsilon\log\rho_{\text{eq}}(m)\to U(m,y)\\ &=\frac{1}{b^{(0)}}\left(\beta_{1}y^{2}\mu+\frac{\beta_{2}y}{2}\mu^{2}+\frac{\sigma}{3}\mu^{3}\right).\end{split}$ (89)
@@
where we recall $\mu=m-m_{c}$ and we define
$\alpha_{1}=\frac{\partial a}{\partial x},\quad\alpha_{2}=\frac{\partial^{2}a}{\partial m\partial y},\quad\sigma=\frac{1}{6}\frac{\partial^{3}a}{\partial m^{3}},$ (93)
@@
For a transverse perturbation to the cusp, we have $x=O(\mu^{3})$ and $y=O(\mu^{3})$, so the second term in Eq. (92) drops out, which implies no bifurcations. For a tangent perturbation to the cusp, we have $x=0$ and $y=O(\mu^{2})$. Thus, by taking into account the leading terms only, we obtain that Eq. (92) becomes
@@
In this case, Eq. (75) agrees with the normal form of pitchfork bifurcation, and we obtain a non-generic 1-parameter family of non-Gaussian distributions near the cusp:
@@ IV.3 Transition time near critical points
$\hat{m}\equiv\frac{m-m_{c}}{\sqrt{-\alpha_{2}y/\sigma}},\quad\hat{x}=\frac{3\sqrt{3}}{2}\cdot\frac{\alpha_{1}\sqrt{\sigma}x}{(-\alpha_{2}y)^{\frac{3}{2}}},\quad\hat{U}=\frac{b_{c}\sigma U}{(\alpha_{2}y)^{2}},$ (102)
such that the rescaled potential is
@@
$\begin{split}&C_{-+}(x,y)=\frac{(\alpha_{2}y)^{2}}{b_{c}\sigma}\hat{C}\left(\frac{3\sqrt{3}}{2}\cdot\frac{\alpha_{1}\sqrt{\sigma}x}{(-\alpha_{2}y)^{\frac{3}{2}}}\right),\\ &C_{+-}(x,y)=C_{-+}(-x,y),\end{split}$ (106)
which are self-similar solutions of the first kind.
@@
Table 1: Summary of scalings of local coordinates and transition rates. The second and third rows apply respectively to Fig. 11 and Fig. 12. The transition rate $C(x)$ describes a transverse perturbation and depends on $x$ only. The transition rate $C(y)$ describes a tangent perturbation and depends on $y$ only. The transition rate $C(x,y)$ describes a general perturbation depending on both $x$ and $y$, where $k$ is a constant and the function $\hat{C}$ is defined in Eq. (105) and plotted in Fig. 13.
Overall, the FP equation works near the two critical curves, although the scalings change when these two curves meet at the cusp; see Table 1 for a summary of scalings of local coordinates and transition rates. The classical IM and the FIM obey the same scaling laws since both belong to the mean-field universality class, but the FIM provides more freedom to control the directions of steady-state bifurcations.
@@ V Discussion
Taken together, our results show that making the pair interaction depend on the system’s own magnetization via linear feedback fundamentally reorders the mean-field phase portrait and the statistics of fluctuations. In contrast with Curie–Weiss—where increasing temperature always erodes bistability and local equilibrium fluctuations are Gaussian—the feedback model exhibits a unit-temperature transcritical point with algebraic relaxation, a re-entrant bistable window that persists up to a finite-temperature fold, and—under small positive field—a cusp so fixed-field slices can display one, two, or even three distinct critical temperatures. Within the wedge of bistability, a Maxwell curve separates the dominance regions; in the feedback case ($\gamma>0$) it rises to the cusp and yields a unique Maxwell temperature for each fixed field in a finite interval, with field increases favoring the upper-magnetization phase and temperature increases favoring the lower one. Near critical sets, a Fokker–Planck reduction of the master equation produces explicit non-Gaussian equilibrium families and barrier functionals that control rare inter-well transitions (exponential lifetimes in system size). Finally, at zero temperature and sufficiently strong feedback, a super-stable mixed phase appears and the full distribution collapses in finite time, providing an exactly solvable instance of finite-time adsorption in a mean-field spin system.
@@
Because a macroscopic observable feedbacks onto the microscopic coupling, the same mechanism offers a compact modeling language for systems where interactions co-evolve with context. In information ecosystems, polarization and platform amplification map naturally to this framework; large-scale evidence on faster and farther diffusion of false news and on echo-chamber formation is consistent with re-entrant polarization, heavy-tailed fluctuations near critical sets, and asymmetric narrative lifetimes predicted by feedback Ising dynamics Vosoughi et al. (2018); Cinelli et al. (2021). In human–AI ecosystems, mean-field mixtures of human and model agents with pair and higher-order interactions display abrupt regime changes as composition and interaction strengths vary, making the Maxwell frontier and barrier-based lifetime scalings useful for alignment/governance diagnostics and for inverse modeling that fits feedback laws to equal-likelihood frontiers and dwell times P. et al. (2022). In center–periphery (empires–provinces) settings Vakulenko and Sudakow (2025); Sudakow et al. (2025), allegiance plays the role of the order parameter while information and resource loops supply feedback; the coexistence of folds and a cusp implies multiple cohesion thresholds, and the Maxwell curve provides a data-driven boundary between center-loyal and provincial-dissent phases—estimable from transition statistics Castellano et al. (2009). Methodologically, natural extensions include spatial/network generalizations (robustness of the wedge and cusp; topology-shifted Maxwell sets), nonlinear or delayed feedbacks (possible oscillatory regimes), and rare-event computation to recover optimal escape paths, building directly on the barrier formulas and normal forms developed here.
@@ Appendix A Derivation and solution of the Fokker-Planck equation for the Curie-Weiss model at the cusp
$\begin{split}&W_{+}(m)=(1+m+2\epsilon)[1-\tanh(m+2\epsilon)]\\ &W_{-}(m)=(1-m+2\epsilon)[1-\tanh(m-2\epsilon)]\\ &W_{0}(m)=1-m\tanh m\end{split}$ (109)

### v2

@@ III.1 Zero external magnetic field
When the temperature approaches zero, $\beta\to\infty$, the $M_{+}$ phase approaches to the plus ground state. A more precise asymptotic
@@ III.2 Nonzero external magnetic field
when $\gamma\to 0$. In the complementary $\gamma\to+\infty$ limit
@@ III.3 Maxwell curve
for the FIM, with ‘potential energy’
$\begin{split}V(m)&=\int_{0}^{m}du\,\text{arctanh}(u)\\ &=m\cdot\text{arctanh}(m)+\frac{1}{2}\log(1-m^{2})\\ &=\frac{(1+m)\log(1+m)+(1-m)\log(1-m)}{2}\end{split}$ (61)
@@
Although a unique Maxwell magnetic field exists at a fixed temperature, a unique Maxwell temperature may not exist at a fixed magnetic field. In the classical IM with $\gamma=0$, we have $h_{*}(T)=0$ as shown in Fig. 3, so there are infinitely many Maxwell temperatures at $h=0$ and none otherwise. Thus, increasing the temperature favors neither the upper nor the lower equilibrium. However, in the FIM with $\gamma>0$, $h_{*}(T)$ increases to the cusp $(T_{c},h_{c})$ as shown in Figs. 33, so there is a unique Maxwell temperature for any fixed $h\in(h_{*}(0),h_{c})$ and none otherwise. For any fixed $h\in(0,h_{c})$, the Maxwell temperature exists in the finite-temperature interval disconnected from $T=0$. The unique feature of the FIM is that the Maxwell temperature is well defined when $h$ is small and either positive or negative, and increasing the temperature favors the lower equilibrium.
@@ III.5 Convergence to stable equilibrium
Figure 10: The evolution of the probability $P_{M_{1}}(t)$ of being in the absorbing state. We chose $M_{0}=2M_{1}+1$. In this case, the adsorption probability is (67). For the three shown curves, $M_{1}+1=30,100,300$. All curves intersect at $t=\log 2$ where $P_{M_{1}}=\frac{1}{2}$. The front around $t=\log 2$ becomes sharper as $M_{1}$ increases.
@@ IV.1 Fokker-Planck equation and approximation error
Hereinafter, we use the variable $m\equiv 2x-1$, and shortly write $\rho$ instead of $\rho(m,t)$. The FP equation in this case reads
@@
$\frac{\log(r)}{2}-\frac{r-1}{r+1}=\frac{(r-1)^{3}}{24}-\frac{(r-1)^{4}}{16}+\ldots$ (78)
@@ IV.2 Equilibrium distributions near critical points
$\alpha=\frac{\partial a}{\partial x},\quad\beta_{1}=\frac{1}{2}\frac{\partial^{2}a}{\partial y^{2}},\quad\beta_{2}=\frac{\partial^{2}a}{\partial m\partial y},\quad\sigma=\frac{1}{2}\frac{\partial^{2}a}{\partial m^{2}},$
@@
$\begin{split}-\epsilon\log\rho_{\text{eq}}(m)&\to U(m,y)\\ &=\frac{1}{b^{(0)}}\left(\beta_{1}y^{2}\mu+\frac{\beta_{2}y}{2}\mu^{2}+\frac{\sigma}{3}\mu^{3}\right).\end{split}$ (88)
@@
where $\mu=m-m_{c}$ and
$\alpha_{1}=\frac{\partial a}{\partial x},\quad\alpha_{2}=\frac{\partial^{2}a}{\partial m\partial y},\quad\sigma=\frac{1}{6}\frac{\partial^{3}a}{\partial m^{3}},$
@@
For a transverse perturbation to the cusp, $x=O(\mu^{3})$ and $y=O(\mu^{3})$, so the second term in Eq. (91) drops out, which implies no bifurcations. For a tangent perturbation to the cusp, $x=0$ and $y=O(\mu^{2})$, so Eq. (91) becomes
@@
In this case, Eq. (75) agrees with the normal form of pitchfork bifurcation. Thus, a non-generic 1-parameter family of non-Gaussian distributions near the cusp reads
@@ IV.3 Transition time near critical points
$\hat{m}\equiv\frac{m-m_{c}}{\sqrt{-\alpha_{2}y/\sigma}},\qquad\hat{x}=\frac{3\sqrt{3}}{2}\cdot\frac{\alpha_{1}\sqrt{\sigma}x}{(-\alpha_{2}y)^{\frac{3}{2}}}$ (100)
and also rescale the potential
$\hat{U}=\frac{b_{c}\sigma U}{(\alpha_{2}y)^{2}}\,.$ (101)
In terms of the rescaled variables, the rescaled potential reads
@@
[…]
effectively depending only on $\hat{x}$ defined in Eq. (100).
@@
Table 1: Summary of scalings of local coordinates and transition rates. The second and third rows apply respectively to Fig. 11 and Fig. 12. The transition rate $C(x)$ describes a transverse perturbation and depends on $x$ only. The transition rate $C(y)$ describes a tangent perturbation and depends on $y$ only. The transition rate $C(x,y)$ describes a general perturbation depending on both $x$ and $y$, where $k$ is a constant, and $\hat{C}$ is defined in Eq. (104) and plotted in Fig. 13.
Overall, the FP equation works near the two critical curves, although the scalings change when these two curves meet at the cusp; see Table 1 for a summary of scalings of local coordinates and transition rates. The classical IM and the FIM obey the same scaling laws, since both belong to the mean-field universality class, but the FIM provides more freedom to control the directions of steady-state bifurcations.
@@ V Discussion
Taken together, our results show that making the pairwise interaction of the IM depend on the system’s own magnetization via a linear feedback function fundamentally reorders the mean-field phase diagram and the statistics of fluctuations.
In contrast with the Curie-Weiss model, where increasing the temperature always erodes bistability, increasing the temperature can favor bistability in the FIM. This temperature-induced bistability endows the FIM with a unit-temperature transcritical bifurcation point at zero magnetic field. Moreover, at small positive magnetic fields, the phase diagram of the FIM exhibits a re-entrant bistable window with a cusp. Therefore, fixed-field slices can display one, two, or even three critical temperatures. Within the wedge of bistability, a Maxwell curve rises to the cusp and separates the dominance regions of the two phases. This yields a unique Maxwell temperature for each fixed field in a finite interval, with field increases favoring the upper-magnetization phase and temperature increases favoring the lower one.
In both the Curie-Weiss model and the FIM, the FP equation describes the dynamics near an equilibrium. Near critical points, this effective description yields explicit families of non-Gaussian equilibrium distributions that relate to the normal forms of the associated steady-state bifurcations. Moreover, these distributions can predict barrier functions and transition rates for rare inter-well transitions with lifetimes exponential in system size. In these settings, the FIM offers more freedom than the Curie-Weiss model to control the directions of steady-state bifurcations due to temperature-induced bistability.
We also note that the FIM exhibits a dynamical finite-size effect at zero temperature: for sufficiently strong feedback, a super-stable MP appears and an initial probability distribution collapses to a Dirac delta function in finite time. As shown in Ref. Ma et al. (2025), super-stability is absent in the Curie-Weiss model unless the coupling is antiferromagnetic. Thus, the FIM provides a minimal model of finite-time collapse of the probability distribution in an exactly solvable mean-field spin system.
The feedback coupling in the FIM can be generalized to an arbitrary function $f(m)$. At zero temperature, one can infer a feedback function $f$ and thus construct a FIM for any bifurcation diagram representing a binary phase transformation Ma et al. (2025). Thus, FIMs can predict the fate of such bifurcation diagrams at finite temperatures without additional information. A key question is how the function $f$ controls key features of the phase diagram at finite temperatures, including critical curves, cusps, and Maxwell curves. Practically, delayed feedback can also be relevant, which may yield oscillatory regimes.
A major application of stochastic bistable systems resembling FIMs is to model rare events defined as transitions between the two stable states. Here, a major goal is to recover optimal escape paths for such transitions, also known as instantons. The classical Freidlin-Wentzell theory calculates the optimal path via the Hamiltonian dynamics of an auxiliary dynamical system Freidlin and Wentzell (2012). A well-known application of this theory is to computing optimal paths in chemical kinetics Dykman et al. (1994). Later, efficient algorithms, e.g., the string method, were developed and widely used to study rare events E et al. (2002). Such methods can be readily applied to calculate optimal paths in FIMs. Near critical curves, our results in Table 1 generalize the classical result that the transition rate exponent is $3/2$ near saddle-nodes Dykman et al. (1994). Thus, the optimal paths near critical curves may inherit some universality from the normal forms. In the low temperature limit, the possible super-stability of the two stable states and the super-instability of the intermediate state may yield key insights into extremely rare events.
The statistical mechanics and dynamics of systems with long-range interactions have attracted much recent interest Campa et al. (2009); Defenu et al. (2023). In this setting, a natural way to construct finite-dimensional FIMs is to generalize the long-range Ising model (LRIM) pioneered by Dyson Dyson (1969), yielding what we call feedback LRIM. Such models contain $p$-spin interactions of the form $J_{p}\propto D_{p}^{-a_{p}}$, $p\geq 2$, where $D_{p}$ is a $p$-spin distance and $a_{p}>0$. Thus, the feedback LRIM approach interpolates between mean-field FIMs for small $a_{p}$’s and nearest-neighbor FIMs for large $a_{p}$’s. The classical LRIM contains only the 2-spin interaction $J_{2}$ with $D_{2}$ being the geodesic distance. A classical renormalization group calculation reveals the location $a_{2}=a_{2}^{*}$ where the critical exponents cross over from long-range to short-range Sak (1973). The equilibrium properties of the classical LRIM near this crossover are still actively explored Angelini et al. (2014); Behan et al. (2017). In feedback LRIMs, new types of crossovers are possible depending on the network topology and also the distance function $D_{p}$, $p\geq 3$, yielding a rich subject to explore further.
The theory of tipping points in open systems applies to climate change, among other fields of recent interest. There are three tipping mechanisms: bifurcation, noise-induced, and rate-dependent Ashwin et al. (2012). At zero temperature, FIMs in a time-varying magnetic field provide minimal models for bifurcation and rate-dependent tipping Ma et al. (2025). For FIMs in the bistable regime at finite temperatures, transitions between the two stable states can be identified as noise-induced tipping.
@@
Because a macroscopic observable feedbacks onto the microscopic coupling, the same mechanism offers a compact modeling language for systems where interactions co-evolve with context. In information ecosystems, polarization and platform amplification map naturally to this framework; large-scale evidence on faster and farther diffusion of false news and on echo-chamber formation is consistent with re-entrant polarization, heavy-tailed fluctuations near critical sets, and asymmetric narrative lifetimes predicted by feedback Ising dynamics Vosoughi et al. (2018); Cinelli et al. (2021). In human–AI ecosystems, mean-field mixtures of human and model agents with pair and higher-order interactions display abrupt regime changes as composition and interaction strengths vary, making the Maxwell frontier and barrier-based lifetime scalings useful for alignment/governance diagnostics and for inverse modeling that fits feedback laws to equal-likelihood frontiers and dwell times Contucci et al. (2022). In center–periphery (empires–provinces) settings Vakulenko and Sudakow (2025); Sudakow et al. (2025), allegiance plays the role of the order parameter while information and resource loops supply feedback; the coexistence of folds and a cusp implies multiple cohesion thresholds, and the Maxwell curve provides a data-driven boundary between center-loyal and provincial-dissent phases—estimable from transition statistics Castellano et al. (2009).
Temperature-induced bistability in FIMs represents a counterintuitive type of noise-induced tipping. Typically, the transition frequency increases as the temperature (noise) increases until bistability is destroyed. However, temperature-induced bistability in FIMs allows the transition frequency to decrease as the temperature increases. Thus, our mean-field model exhibits high-temperature ordered phases in a restricted sense. This defies conventional wisdom. Interestingly, ordered phases persisting in arbitrarily high temperatures have been recently reported Han et al. (2025); Komargodski and Popov (2025); Huang et al. (2025), and named entropic order, in quantum and classical many-body systems.
@@ Appendix A Derivation and solution of the Fokker-Planck equation for the Curie-Weiss model at the cusp
$\begin{split}&W_{+}(m)=(1+m+2\epsilon)[1-\tanh(m+2\epsilon)]\\ &W_{-}(m)=(1-m+2\epsilon)[1+\tanh(m-2\epsilon)]\\ &W_{0}(m)=1-m\tanh m\end{split}$ (108)